OpenStudy (anonymous):

Five percent of a number is 11 more than 3% of the number. Find the number.

OpenStudy (anonymous):

\[0.05n = 11 + 0.03n\] \[0.02n = 11\] \[n = 550\]

OpenStudy (anonymous):

would you help me with another

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What percent of 7.6 pounds is 30 pounds?

OpenStudy (anonymous):

\[\frac{ 30 }{ 7.6 }*100 = 394.7\]